# A Flexible Plug-and-Play Module for Generating Variable-Length

### v1

@@ 3.4. Adaptive Weights Strategy
(5) $\alpha_{i}cos\theta_{ik}^{(k)}\lVert g_{i}^{(k)}\rVert+\frac{\alpha_{k}}{m-k}\lVert g_{k}^{(k)}\rVert\geq 0;\ k\leq i\leq m.$
It is easy to demonstrate that if Eq. (5) holds for any $k$ and $i$, then Eq. (4) also holds. Through Eq.(5), we can get the range of $\alpha_{i}$. Without loss of generality, we first set $\alpha_{1}=1$ for the shortest code’ objective $\mathcal{L}_{1}$, as normalization can subsequently be applied. Then we introduce $\alpha_{i}^{(k)}$ denote only consider to ensure that $L_{i}$ and $L_{k}$ satisfy Eq. (5) on $W^{(k)}$. Using $cos\theta_{ik}^{(k)}=\frac{g_{i}^{(k)}g_{k}^{(k)}}{\lVert g_{i}^{(k}\rVert\lVert g_{k}^{(k)}\rVert}$ and re-arranging terms, we then get:
(6) $\alpha_{i}^{(k)}(-g_{i}^{(k)}g_{k}^{(k)})\leq\frac{\alpha_{k}}{m-k}\lVert g_{k}^{(k)}\rVert^{2};\ k\leq i\leq m.$
(7) $\alpha_{i}^{(k)}\leq\frac{\alpha_{k}}{k-m}\frac{\lVert g_{k}^{(k)}\rVert^{2}}{g_{i}^{(k)}g_{k}^{(k)}};\ k\leq i\leq m.$
(8) $\alpha_{i}^{(k)}=\left\{\begin{array}[]{rcl}1&\text{if}&g_{i}^{(k)}\cdot g_{k}^{(k)}\geq 0\\ \frac{\alpha_{k}}{k-m}\frac{\lVert g_{k}^{(k)}\rVert^{2}}{g_{i}^{(k)}g_{k}^{(k)}}&\text{if}&g_{i}^{(k)}\cdot g_{k}^{(k)}<0;k\leq i\leq m\\ \end{array}\right.$
(9) $\alpha_{i}=min(\alpha_{i}^{(1)},\alpha_{i}^{(2)},...,\alpha_{i}^{(i)}).$
In each training step, we dynamically compute the $\{\alpha_{k}\}_{k=1}^{m}$ using Eq. (8) and Eq. (9). The computation complex is $O(lb_{m}m^{2})$, where $l$ is the dimension of data feature $v$ and $b_{m}$ is the longest code length. We conducted an experiment in Section 5.5 to demonstrate that incorporating the adaptive weights strategy increases the training time per step by approximately 11.15%.
3.5. Long-short Cascade Self-distillation
Compared to the traditional hash layer, the deep hashing model with NHL can simultaneously generate hash codes with varying code lengths. This prompts us to consider the intrinsic connections among hash codes of varying lengths. We find a teacher-student relationship is established between long hash codes and short hash codes. Therefore, As shown in Figure 2 (c), we propose the long-short cascade self-distillation method, leveraging long hash codes to enhance the performance of short hash codes in a cascade manner.
@@ 3.5. Long-short Cascade Self-distillation
(10) $\mathcal{L}_{k}^{lcs}=\frac{1}{B}\lVert\frac{h_{i}^{(k)}H_{k}^{T}}{\lVert h_{i}^{(k)}H_{k}^{T}\rVert}-\frac{h_{i}^{(k+1)}H_{(k+1)}^{T}}{\lVert h_{i}^{(k+1)}H_{(k+1)}^{T}\rVert}\rVert^{2}.$
@@
(11) $\mathcal{L}=\sum_{i=1}^{N}\sum_{k=1}^{m-1}\alpha_{k}(\mathcal{L}_{k}+\lambda\mathcal{L}_{k}^{lcs})+\sum_{i=1}^{N}\alpha_{m}\mathcal{L}_{m},$
@@
We renormalize the weights $\alpha_{k}$ in each step so that $\sum_{k=1}^{m}\alpha_{k}=m$ to decouple gradient re-weight from the global learning rate. Besides, in the training procedure, note that the minimum of $\mathcal{L}$ does not necessarily imply that each $\{\mathcal{L}_{k}\}_{k=1}^{m}$ is at its minimal value during the training process. Therefore, we propose a trick for our training procedure. Throughout the training, we monitor the value of each $\mathcal{L}_{k}$ and save the model parameters when each $\mathcal{L}_{k}$ reaches its minimum to output the corresponding hash codes $h^{(k)}$. We summarize the whole algorithm in Algorithm 1. We present the training algorithm of our proposed NHL in Algorithm 1. In lines 12-14 of Algorithm 1, when a smaller $L_{k}$ is achieved, We record the current model parameters $\theta_{F}$ and $W$ as the parameters of the model with a length of $b_{k}$, denoted as $\theta_{F}^{(k)}$ and $W^{(k)}$.
Algorithm 1 The training algorithm with NHL
1:  Initialization: the parameter of the deep hashing model $\{\theta_{F},W\}$, $\alpha_{k}=1\ \forall k$.
2:  repeat
4:     for each $k\in\{1,2,....,m\}$ do
5:        for each $i\in\{1,2,....,m\}$ do
6:           obtain $g_{i}^{(k)}$ by computing standard gradients $g_{i}^{(k)}=\frac{\partial\mathcal{L}_{i}}{\partial W^{(k)}}$Only calculating the gradients on $\{W^{(k)}\}_{k=1}^{m}$}
7:        end for
8:     end for
9:     compute $\{\alpha_{k}\}_{k=1}^{m}$ by Eq. (9) and Eq. (8)
10:     renormalize $\{\alpha_{k}\}_{k=1}^{m}$ so that $\sum_{k=1}^{m}\alpha_{k}=m$
12:     if achieved a smaller $\mathcal{L}_{k}$ then
13:        record the current model parameters $\theta_{F}^{(k)}$, $W^{(k)}$ for the model of length $b_{k}$.
14:     end if
15:  until converged
15:  parameters of deep hashing model $\{\theta_{F}^{(k)}\}_{k=1}^{m}$ and $\{W^{(k)}\}_{k=1}^{m}$
4. Discussion & Future Direction
The proposed NHL module can be integrated into the majority of deep supervised hashing models (Liu et al., 2016; Zhu et al., 2016, 2017; Cao et al., 2018; Li et al., 2020; Zheng et al., 2020; Wang et al., 2017; Liu et al., 2018; Cakir et al., 2019; Yuan et al., 2020; Fan et al., 2020; Hoe et al., 2021; Wang et al., 2022, 2023; He et al., 2024b). A characteristic feature of these models is that both the database data and the query data rely on the same deep hashing network for hash code generation, known as symmetric method. Nevertheless, we must clarify that NHL is not applicable to some deep supervised hashing models (Shen et al., 2017; Jiang and Li, 2018; Chen et al., 2019; Wu et al., 2023), which are known as the two-step method or asymmetric method. These models are characterized by their direct or indirect optimization of hash codes in the database rather than obtaining hash codes through the output of the deep neural network. They solely utilize deep neural networks for processing query images and obtaining their hash codes. Exploring how to adapt NHL to these models represents a worthwhile direction for future research.
Besides, in our earlier experiments, we also explored applying NHL in deep unsupervised hashing models. However, we found that the NHL does not consistently achieve significant improvements in these models. This underscores the importance of supervised signals in multiple objective learning. Exploring how to integrate the concept of NHL into deep unsupervised hashing models remains a worthwhile pursuit.
5. Experiments
5.1. Experiment Settings
5.1.1. Dataset
We conducted experiments on three widely used datasets in deep hashing for evaluation. CIFAR-10 (Krizhevsky et al., 2009) consists of 60,000 images from 10 classes. Following (Cao et al., 2018), we randomly select 1,000 images per class as the query set, and 500 images per class as the training set, and use all remaining images as the database. ImageNet100 is a subset of ImageNet (Deng et al., 2009) with 100 classes. We follow the settings from (Fan et al., 2020) and randomly select 100 categories. Then, we use all the images of these categories in the training set as the database and the images in the validation set as the queries. Furthermore, we randomly select 13,000 as the training images from the database. MSCOCO (Lin et al., 2014) is a large-scale dataset for object detection, segmentation, and captioning. We consider a subset of 122,218 images from 80 categories, as in previous works (Qiu et al., 2021). We randomly select 5,000 images from the subset as the query set, and use the remaining images as the database. For training, we randomly select 10,000 images from the database. As in most deep hashing settings, two samples are viewed as similar if they correspond to the same label on CIFAR-10 and ImageNet100. For multi-label datasets MSCOCO, two samples are considered similar if they share at least one common label.
Data Model w/o NHL w/ NHL
@@ 5.1.1. Dataset
CIFAR-10 (mAP@1000) DSH 0.690 0.731 0.740 0.727 0.381 0.654 $\boldsymbol{0.717}$ 0.732 0.744 $\boldsymbol{0.743}$ $\boldsymbol{0.749}$ $\boldsymbol{0.737}$ $(+12.8\%)$
@@
Table 1. The mAP@K comparison results on CIFAR-10, ImageNet100, and MSCOCO datasets when different deep hashing models used the original hash layer (w/o NHL) or NHL (w/ NHL). We employ bold numbers to indicate statistically significant enhancements when utilizing NHL compared to when not using NHL, with $p<0.05$ based on a two-tailed paired t-test.
5.2. Evaluation Metric
We employed the mean Average Precision at the top K ($mAP@K$) as the evaluation metric. Specifically, we utilized $mAP@5000$ for MSCOCO and $mAP@1000$ for both CIFAR-10 and ImageNet100, following the settings used in previous studies (Qiu et al., 2021; Fan et al., 2020). Unless otherwise specified, we set the hash code lengths $b\in\{8,16,32,64,128\}$ for the following experiments, as these lengths are prevalently used in previous works.
5.3. Baselines and Training Details
We considered the following deep supervised hashing models, including DSH (Liu et al., 2016), DTSH (Wang et al., 2017), LCDSH (Zhu et al., 2017), DCH (Cao et al., 2018), CSQ (Yuan et al., 2020), DPN (Fan et al., 2020), and MDSH (Wang et al., 2023). In subsequent experiments, we use w/o NHL to represent deep hashing models without using NHL, and w/ NHL means deep hashing models use NHL to replace the original hash layer. We carefully implemented all models using Pytorch and conducted experiments on an NVIDIA Geforce RTX4090 GPU. We used ResNet50 (He et al., 2016) as the backbone to extract 2048-D image features and try our best to re-implement the previous methods. The batch size $B$ was set to $64$. When deep hashing models applied NHL, we used the Adam optimizer (Kingma and Ba, 2014) and selected the learning rate from $\{10^{-4},10^{-5}\}$. The hyper-parameter $\lambda$ was selected from $\{10^{1},10^{0},10^{-1},10^{-2},10^{-3}\}$. We perform the grid search method on different cases for the best combination.
5.4. Performance on Deep Hashing Models
In this experiment, we compared the $mAP@K$ of different deep supervised hashing models on three datasets. Table 1 presents an overview of the results. We use “w/o NHL” to denote the deep hashing model without using NHL and use “w/ NHL” to denote the deep hashing model that uses NHL to replace the original hash layer. In its original paper, MDSH (Wang et al., 2023) does not present methods for processing multi-label datasets. Thus, we cannot get the results of MDSH on the MSCOCO dataset.
@@ 5.4. Performance on Deep Hashing Models
Table 1 shows the results. We use bold numbers to indicate statistically significant improvements when utilizing NHL compared to not using NHL, with $p<0.05$ based on a two-tailed paired t-test. We can find the following observations: (i) Globally, the implementation of the NHL leads to an average improvement of 4.46% (4.69% in CIFAR-10, 6.03% in ImageNet100, and 2.35% in MSCOCO). Besides, there are 72% of cases that achieve a significant performance boost based on the two-tailed paired t-test. Conversely, only a few cases achieve a decline, with most drops of 1.37% occurring in the DTSH model when NHL is applied to the ImageNet100 dataset using a 128-bit code. Thus, we can demonstrate that NHL can yield significant improvements in the majority of cases. (ii) Deep hashing models with NHL improve significantly when the hash code length is short in some datasets. For example, in the case of 8-bit, employing NHL can increase 18.7% and 7.32% enhancement on ImageNet100 and MSCOCO datasets, respectively. This is attributed to the adaptive weights strategy and long-short cascade self-distillation introduced for the NHL structure, ensuring optimization for short hash codes and leveraging the relationships in long hash codes to enhance the performance of short hash codes. (iii) It is delightful to note that NHL can address the curse of dimensionality of hash code, signifying that as the code length expands to a certain dimension, the code quality commences to deteriorate in some deep hashing models. For example, without NHL, the quality of hash codes in DSH experiences a marked decline when transitioning from 64 bits to 128 bits on the CIFAR-10 dataset. In contrast, with the incorporation of the NHL, this result undergoes a substantial improvement.
@@
Table 2. The efficiency evaluation on three datasets. We record the total training time for the deep hashing model of five code lengths and the maximal memory usage.
5.5. Efficiency Analysis
In this experiment, we evaluated the deep hashing model’s training time and memory usage. We recorded the total training time for the hashing model of five code lengths and recorded the maximal memory usage.
Table 2 displays the results, where we selected CSQ and DCH as the deep hashing models. These results demonstrate that the employment of the NHL incurs negligible additional memory expenses. This is attributed to the fact that during the training process, the primary memory usage stems from the parameters of the neural network, while the additional memory occupied by the target loss is relatively minimal. Meanwhile, the incorporation of the NHL can significantly enhance the overall training speed. Specifically, it achieved an average training speedup of 5.94×5.94$\times$5.94 ×, 6×6$\times$6 ×, and 5.31×5.31$\times$5.31 × on the CIFAR-10, ImageNet-100, and MSCOCO datasets, respectively, across the two deep hashing models. In conjunction with the conclusions drawn from the previous experiment, this evidences that NHL can effectively expedite the training procedure without compromising the quality of hash codes.
Besides, we further analyze the average time per epoch during training across different NHL variants. Figure 4 displays the results of CSQ and DCH. Compared to deep hashing models without NHL (w/o NHL), the use of NHL-basic and NHL w/o A resulted in an increase in training time of 3.37% and 6.87%, respectively. This indicates that introducing a basic NHL structure and employing self-distillation do not significantly extend the duration of each epoch. Furthermore, compared to models without using NHL (w/o NHL), those with NHL w/o D and w/ NHL experience an average increase in training time of 11.15% and 13.75%, respectively. This increase is primarily attributed to our adaptive strategy that explicitly monitors and analyzes parameters within NHL rather than across the entire model.
5.6. Ablation Study
To analyze the influence of each component in the NHL, we conducted an ablation study on these models to investigate their impact. We devised several variants for the NHL, namely (i) NHL-basic: directly use E.q (2) to optimize the deep hashing model, (ii) NHL w/o A: without using the adaptive weights strategy, (iii) NHL w/o D: without using the long-short cascade self-distillation. We present the results CSQ on CIFAR-10 and ImageNet100 in Figure 5. We can find that employing NHL-basic does not enhance the performance of hash codes compared to using the original hash layer. This indicates that simply jointly learning the objective function like Eq.(2) is insufficient. Furthermore, the adaptive weight strategy (NHL w/o D) and the long-short cascade self-distillation (NHL w/o A) enhance the performance of NHL-basic. These results underscore the significance of contemplating gradient optimization scenarios and facilitating short hash codes to learn from long code relationships. Finally, a notable enhancement can be achieved by simultaneously considering these two factors (w/ NHL).
(a) DCH
(b) DTSH
@@ 5.6. Ablation Study
Figure 4. The average training time per epoch for CSQ and DCH on three datasets, when they use the original hash layer, NHL, or the variants of NHL.
(a) CIFAR-10
(b) ImageNet100
Figure 5. The $mAP@K$ results with NHL variants on CIFAR-10 and ImageNet100 datasets.
5.7. Module Analysis
In this experiment, we conducted a comprehensive analysis of NHL from multiple perspectives, including (i) The adaptation to transformer backbones, (ii) hyper-parameter analysis, (iii) gradient analysis, and (iv) More code length settings. We use CSQ as the deep hashing model in the following analysis.
5.7.1. Adaptation to Transformer Backbones
The proposed NHL module can be integrated into the majority of deep supervised hashing models. In most deep hashing models, deep convolutional neural networks like ResNet50 (He et al., 2016) are used to extract the image feature, and ResNet50 also served as the backbone for our prior experiments. Recently, the transformer architecture has achieved success in various fields. Some studies (Gong et al., 2022; He et al., 2024a) on deep hashing have explored the ViT (Dosovitskiy et al., 2020) to extract the image feature and get better results. To evaluate the compatibility of the NHL with the transformer backbone, we used ViT_B_16 as the backbone of the CSQ model. Figure 6 shows the results, revealing that the NHL remains effective when integrated with the ViT_B_16 backbone, signifying that our proposed NHL is equally applicable to the current transformer architectures.

### v2

@@ 3.4 Dominance-Aware Dynamic Weighting strategy
$\alpha_{k}\lVert g_{k}^{(k)}\rVert+\sum_{k<i\leq m}\alpha_{i}cos\theta_{ik}^{(k)}\lVert g_{i}^{(k)}\rVert\geq 0,$ (4)
then the update of $\mathcal{W}^{(k)}$ is guaranteed to be align-domination.
The proof is provided in Appendix B. However, the linear programming Eq.(4) is challenging to optimize and will incur additional time expenditure. Hence, we propose the following target and proposition:
Proposition 2.
If the following inequality holds for all $k<i\leq m$, then Eq. (4) also holds.
$\alpha_{i}cos\theta_{ik}^{(k)}\lVert g_{i}^{(k)}\rVert+\frac{\alpha_{k}}{m-k}\lVert g_{k}^{(k)}\rVert\geq 0.$ (5)
We provide the proof in Appendix B and introduce a method to solve Eq. (5) in Appendix C. Here, for arbitrary $k\leq i\leq m$, we directly provide the results:
$\alpha_{i}=min(\alpha_{i}^{(1)},\alpha_{i}^{(2)},...,\alpha_{i}^{(i)}).$ (6)
$\alpha_{i}^{(k)}=\left\{\begin{array}[]{rcl}1&\text{if}&g_{i}^{(k)}\cdot g_{k}^{(k)}\geq 0\\ \frac{\alpha_{k}}{k-m}\frac{\lVert g_{k}^{(k)}\rVert^{2}}{g_{i}^{(k)}g_{k}^{(k)}}&\text{if}&g_{i}^{(k)}\cdot g_{k}^{(k)}<0.\\ \end{array}\right.$ (7)
In each training step, we dynamically compute the $\{\alpha_{k}\}_{k=1}^{m}$ using Eq. (6) and Eq. (7). Similar to chen2018gradnorm , we don’t consider the full network weights and focus on the parameter in NHL. The computation complex is $O(lb_{m}m^{2})$, where $l$ is the dimension of data feature $v$ and $b_{m}$ is the longest code length. Our experiment shows that the additional training time for each step is around 11.15%.
3.5 Long-short Cascade Self-distillation
Unlike the traditional hash layer, the deep hashing model with NHL can generate hash codes of different lengths simultaneously. This leads us to explore the connections among these hash codes of different lengths. We observe a teacher-student relationship between long and short hash codes. Thus, as illustrated in Figure 2c, we propose the long-short cascade self-distillation method, which uses long hash codes to improve the performance of short hash codes in a cascading manner.
@@ 3.5 Long-short Cascade Self-distillation
$\mathcal{L}_{k}^{lcs}=\frac{1}{B}\left\lVert\frac{h_{i}^{(k)}H_{k}^{T}}{\lVert h_{i}^{(k)}H_{k}^{T}\rVert}-\frac{h_{i}^{(k+1)}H_{(k+1)}^{T}}{\lVert h_{i}^{(k+1)}H_{(k+1)}^{T}\rVert}\right\rVert^{2}.$ (8)
@@
$\mathcal{L}=\sum_{i=1}^{N}\sum_{k=1}^{m-1}\alpha_{k}(\mathcal{L}_{k}+\lambda\mathcal{L}_{k}^{lcs})+\sum_{i=1}^{N}\alpha_{m}\mathcal{L}_{m},$ (9)
@@
We renormalize the weights $\alpha_{k}$ in each step so that $\sum_{k=1}^{m}\alpha_{k}=m$ to decouple gradient re-weight from the global learning rate. Besides, in the training procedure, note that the minimum of $\mathcal{L}$ does not necessarily imply that each $\{\mathcal{L}_{k}\}_{k=1}^{m}$ is at its minimal value during the training process. Therefore, we propose a trick for our training procedure. Throughout the training, we monitor the value of each $\mathcal{L}_{k}$ and save the model parameters when each $\mathcal{L}_{k}$ reaches its minimum to output the corresponding hash codes $h^{(k)}$. We summarize the whole algorithm in Appendix D.
4 Experiments
4.1 Experiment Settings
We evaluated our method on three widely used datasets in deep hashing: CIFAR-10 krizhevsky2009learning , ImageNet100 deng2009ImageNet , and MSCOCO lin2014microsoft . We compared our approach against several state-of-the-art deep supervised hashing baselines: DSH liu2016deep , DHN zhu2016deep , DTSH wang2017deep , LCDSH zhu2017locality , DCH cao2018deep , DBDH zheng2020deep , CSQ yuan2020central , SHCIR wang2022image , DPN fan2020deep , and MDSH wang2023deep . For all the above models, we adopted ResNet50 he2016deep as the backbone. The primary evaluation metric was mean Average Precision at top K ($mAP@K$). Unless otherwise specified, we set the hash code lengths $b\in\{8,16,32,64,128\}$ for the following experiments, as these lengths are prevalently used in previous works. Details regarding datasets, implementation, and evaluation settings are presented in Appendix A.
Table 1: The mAP@K comparison results on CIFAR-10, ImageNet100, and MSCOCO datasets when different deep hashing models used the original hash layer (w/o NHL) or NHL (w/ NHL). We employ bold numbers to indicate statistically significant enhancements when utilizing NHL compared to when not using NHL, with $p<0.05$ based on a two-tailed paired t-test.
Data Model w/o NHL (The Original Model) w/ NHL
@@ 4.1 Experiment Settings
CIFAR-10 (mAP@ALL) DSH 0.690 0.731 0.740 0.727 0.381 0.654 $\boldsymbol{0.717}$ 0.732 0.744 $\boldsymbol{0.743}$ $\boldsymbol{0.749}$ $\boldsymbol{0.737}$ $(+12.8\%)$
@@
SHCIR 0.789 0.861 0.879 0.883 0.881 0.858 $\boldsymbol{0.798}$ $\boldsymbol{0.881}$ $\boldsymbol{0.889}$ $\boldsymbol{0.893}$ $\boldsymbol{0.898}$ $\boldsymbol{0.872}$ $(+1.63\%)$
@@
DBDH 0.655 0.727 0.760 0.769 0.800 0.742 $\boldsymbol{0.692}$ $\boldsymbol{0.748}$ $\boldsymbol{0.778}$ $\boldsymbol{0.803}$ $\boldsymbol{0.809}$ $\boldsymbol{0.766}$ $(+3.23\%)$
@@ 4.2 Performance on Deep Hashing Models
In this experiment, we first compared the $mAP@K$ of different deep supervised hashing models on three datasets. Table 1 shows the results. We use “w/o NHL” to denote the deep hashing model without using NHL and use “w/ NHL” to denote the deep hashing model that uses NHL to replace the traditional hash layer. Besides, we use bold numbers to indicate statistically significant improvements when utilizing NHL compared to not using NHL, with $p<0.05$ based on a two-tailed paired t-test.
@@
We can find the following observations: (i) Globally, the implementation of the NHL leads to an average improvement of 3.398% (3.619% in CIFAR-10, 4.364% in ImageNet100, and 2.158% in MSCOCO). Besides, there are 72% of cases that achieve a significant performance boost based on the two-tailed paired t-test. Conversely, only a few cases achieve a decline, with most drops of 1.37% occurring in the DTSH model when NHL is applied to the ImageNet100 dataset using a 128-bit code. Thus, we can demonstrate that NHL can yield significant improvements in the majority of cases. (ii) Deep hashing models with NHL improve significantly when the hash code length is short in some datasets. For example, in the case of 8-bit, employing NHL can increase 18.7% and 7.32% enhancement on ImageNet100 and MSCOCO datasets, respectively. (iii) It is delightful to note that NHL can address the dimensionality curse of hash code, signifying that as the code length expands to a certain dimension, the code quality commences to deteriorate in some deep hashing models. For example, without NHL, the quality of hash codes in DSH experiences a marked decline when transitioning from 64 bits to 128 bits on the CIFAR-10 dataset. In contrast, with the incorporation of the NHL, this result undergoes a substantial improvement.
Table 2: The comparison of average mAP@K results with the original model is shown for NHL variants.
Data NHL-basic NHL w/o D NHL w/o L w/ NHL
CIFAR-10 +1.088% +1.316% +2.097% +3.619%
ImageNet100 +2.209% +2.823% +2.434% +4.364%
MSCOCO +0.421% +0.859% +0.978% +2.158%
avg. +1.228% +1.639% +1.856% +3.398%
Besides, to analyze the influence of each component in the NHL, we conducted an ablation study on these models to investigate their impact. We devised several variants for the NHL, namely (i) NHL-basic: directly use E.q (2) to optimize the deep hashing model, (ii) NHL w/o D: without using the dominance-aware dynamic weighting strategy, (iii) NHL w/o L: without using the long-short cascade self-distillation. Table 2 presents the average performance changes across datasets for the models mentioned above. Overall, implementing NHL-basic, NHL w/o D, and NHL w/o L results in average performance improvements of 1.228%, 1.639%, and 1.856%, respectively. Employing only the dominance-aware dynamic weighting strategy (NHL w/o L) achieves the most significant relative improvement, highlighting the critical role of gradient optimization in this context. In Appendix F, we present the detailed results of various deep hashing models when utilizing the different variants of NHL.
Table 3: We evaluate efficiency on three datasets by recording the total training time for the deep hashing model with five code lengths and the maximum memory usage.
@@
(a) CSQ
(b) DCH
@@
Figure 4: The average training time per epoch for CSQ and DCH across three datasets, using the traditional hash layer, NHL, or NHL variants.
4.3 Efficiency Analysis
In this experiment, we evaluated the deep hashing model’s training time and memory usage. We recorded the total training time for the hashing model of five code lengths and recorded the maximal memory usage. Table 4 displays the results, where we selected CSQ and DCH as the deep hashing models. These results demonstrate that the employment of the NHL incurs negligible additional memory expenses. This is attributed to the fact that during the training process, the primary memory usage stems from the parameters of the neural network, while the additional memory occupied by the target loss is relatively minimal. Meanwhile, the incorporation of the NHL can significantly enhance the overall training speed. It achieved an average training speedup of 5.94×5.94$\times$5.94 ×, 6×6$\times$6 ×, and 5.31×5.31$\times$5.31 × on the CIFAR-10, ImageNet-100, and MSCOCO datasets, respectively, across the two deep hashing models. In conjunction with the conclusions drawn from the previous experiment, this evidences that NHL can effectively expedite the training procedure without compromising the quality of hash codes.
Besides, we further analyze the average time per epoch during training across different NHL variants. Figure 44(a) and Figure 44(b) display the results of CSQ and DCH. Compared to deep hashing models without NHL (w/o NHL), incorporating NHL-basic, NHL w/o D, NHL w/o L, and w/ NHL led to a modest increase in training time of just 3.37%, 6.87%, 11.15%, and 13.75%, respectively.
4.4 Module Analysis
In this experiment, we conducted a comprehensive analysis of the Nested Hash Layer (NHL) from multiple perspectives, including (i) hyperparameter analysis and (ii) various code length settings. We utilize CSQ as the deep hashing model for the subsequent analysis. Additionally, we validated the performance of the Nested Hash Layer (NHL) under different backbone extractors, including other variants of ResNet and different architectures of Vision Transformers (ViT) dosovitskiy2020image . For further details, please refer to Appendix E.2.
4.4.1 Parameter Sensitivity
In Eq.(9), $\lambda$ serves as a hyperparameter that balances two objectives. We evaluated its values from $\{10^{1},10^{0},10^{-1},10^{-2},10^{-3}\}$ to calculate the $mAP@K$ across three datasets. Figure 6 illustrates the results, highlighting that the performance of NHL demonstrates overall robustness to changes in $\lambda$. Generally, the optimal value is achieved when $\lambda=1$. Additionally, we conducted a parameter analysis on the learning rate, which is presented in Appendix E.1. The analysis indicates that the optimal range for the learning rate is $\{10^{-4},10^{-5}\}$.
4.4.2 More Code Length Settings
This section explores the results under a broader range of code length settings. We established three scenarios for code length: Case 1 sets a code length at every 32-bit interval, that is, $b=\{32\times k\}_{k=1}^{4}$ and $m=4$. Case 2 sets a code length at every 16-bit interval, that is, $b=\{16\times k\}_{k=1}^{8}$ and $m=8$. Case 3 sets a code length at every 8-bit interval, that is, $b=\{8\times k\}_{k=1}^{16}$ and $m=16$. Figure 6 presents the corresponding results. Here, the blue bars represent the average ratio of $mAP@K$ at various lengths with and without using NHL. The yellow bars indicate the time cost ratio to complete training with and without using NHL. We observe that even with different code length settings, the use of NHL ensures a reduction in overall training time and improves the code quality. Moreover, it is noteworthy that the total training time does not monotonically increase with the number of output code lengths. For instance, the efficiency enhancement ratio in Case 3 is not as high as in Case 2. This is attributed to the requirement for the model to undergo more training iterations in Case 3 to ensure favorable outcomes across a greater number of code lengths.